Accelerator Update for December 3, 2008 - December 5, 2008
Wednesday Dec 3
The day shift began with the Collider Detector at Fermilab (CDF) in access, with Operations monitoring the Recycler stash (385.9E10), with the Antiproton Source stacking (14.3E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the D-Zero (D0) experiment waiting for the next store, and with the MiniBooNE experiment taking beam. NuMI remained down to replace their horn.
At 8:14 AM, Operations put Pbar into standby mode while Booster experts investigated a problem with the East Anode Power Supply. Pbar resumed stacking at 8:31 AM.
The Neutron Therapy Facility reported at 9:47 AM that they had completed treatment on their last patient for the day.
Operations reported at 10:14 AM that CDF had come out of access and operators had completed their search and secure of the collision hall.
At 11:51 AM, the TeV abort dropped on a RF LCW temperature alarm indication. Operations contacted an expert who said he had pulled out the temperature monitor to look at it, but it was not connected to the RF abort. Operations had him pull the module again and it triggered another abort. A third test did not cause an abort. The expert thought the crate power supply might have been the cause and would investigate further.
Operations put the TeV into shot setup at 12:28 PM.
Operations established store 6609 at 2:22 PM with an initial luminosity of 314.46E30.
A Pbar expert began a study at 3:19 PM.
A Recycler expert began an Electron Cooling (ECool) study at 3:22 PM.
Operations reported at 4 PM that the Recycler had stashed antiprotons once during the day shift.
The ECool study ended and experts accessed the Pelletron. That access ended at 4:12 PM.
Operations removed the slipstacking event from the Time Line Generator at 5:43 PM.
At 8:05 PM, Operations reported a noise problem with a Linac quadrupole power supply (L:QPS407). Operators adjusted the supply.
At 11:57 PM, the TeV quenched at sectors C4 and D1.
Operations reported at midnight that the Recycler had stashed six times during the evening shift.
Thursday Dec 4
The midnight shift began with Operations monitoring the stash (236.2E10), with Pbar stacking (47.6E10), and with MiniBooNE taking beam.
Cryo experts reported at 12:30 AM that they were investigating the TeV quench. Later they reported that the quench was due to a controls problem.
At 3:14 AM, cryo system experts reported that they were having trouble with the TeV sector C4 cold compressor.
Operations put the TeV through a dry squeeze at 6:02 AM.
At 6:33 AM, the TeV aborted at sector F1 while operators were loading protons. Operations investigated and discovered a power supply (T:QDD1) that was not putting out any current. Operations contacted an expert who came in and began repairs.
Operations reported at 8:01 AM that the Recycler had stashed seven times during the midnight shift.
Pbar experts began a cooling study at 8:57 AM. This halted stacking for several minutes.
Operations put the TeV through a dry squeeze at 9:40 AM.
At 9:44 AM, the beam output of Linac dropped. Experts investigated.
Operations put the TeV through a wet squeeze at 10:12 AM.
Operations put the TeV into shot setup at 11:04 AM.
Operations established store 6614 at 12:44 PM with an initial luminosity of 336.2E30.
Pelletron experts made an access into MI-31 at 1:11 PM. They were out 40 minutes later.
A MI expert began a slip stacking study at 2:43 PM.
Operations reported at 4:02 PM that the Recycler had stashed three times during the day shift.
At 5:44 PM, an expert responded to a report earlier today of a smell of burnt electronic at D0. He found two power supplies with bad SCRs (C:D0Q5 and Q6). The plot on the left shows a good output. The plot on the right shows D0Q6’s output.
At 6:07 PM, Operations reported that D:EKIK was having firing problems. Operations contacted an expert.
Operations terminated store 6614 at 8:13 PM and prepared the D0 collision hall for access.
Power supply experts began their repairs on D0Q5 and 6 at 8:43 PM.
D0 came out of access at 8:58 PM.
Power supply experts completed their work on D0Q5 and 6 at 10:43 PM. They found corrosion on a Q5 SCR firing card and a loose lead on a Q6 SCR firing card.
At 11:17 PM, D:EKIK tripped off and would not reset. Operations called in an expert and put Pbar into standby mode. The expert had the kicker fixed before the end of shift.
Operations put the TeV through a dry squeeze at 11:32 PM.
Operations reported at midnight that the Recycler had stashed seven times during the evening shift.
Friday Dec 5
The midnight shift began with Operations monitoring the stash (305.9E10), with Pbar stacking (39.8E10), and with MiniBooNE taking beam.
Operations put the TeV through a wet squeeze at 12:23 AM.
Operations put the TeV into shot setup at 2:11 AM.
Operations established store 6619 at 3:34 AM with an initial luminosity of 317.95E30.
Operations reported at 8 AM that the Recycler had stashed seven times during the midnight shift.
The Plans for Friday and the Weekend
The plans for today and the weekend are to stack, stash, and store.
Comments and Suggestions
|last modified 01/05/2009 email Fermilab|
|Security, Privacy, Legal|